# VisRL: Intention-Driven Visual Perception via Reinforced Reasoning

### v1

@@ Abstract
Visual understanding is inherently intention-driven—humans selectively focus on different regions of a scene based on their goals. Recent advances in large multimodal models (LMMs) enable flexible expression of such intentions through natural language, allowing queries to guide visual reasoning processes. Frameworks like Visual Chain-of-Thought have demonstrated the benefit of incorporating explicit reasoning steps, where the model predicts a focus region before answering a query. However, existing approaches rely heavily on supervised training with annotated intermediate bounding boxes, which severely limits scalability due to the combinatorial explosion of intention-region pairs. To overcome this limitation, we propose VisRL, the first framework that applies reinforcement learning (RL) to the problem of intention-driven visual perception. VisRL optimizes the entire visual reasoning process using only reward signals. By treating intermediate focus selection as a internal decision optimized through trial-and-error, our method eliminates the need for costly region annotations while aligning more closely with how humans learn to perceive the world. Extensive experiments across multiple benchmarks show that VisRL consistently outperforms strong baselines, demonstrating both its effectiveness and its strong generalization across different LMMs. Our code is available at https://github.com/zhangquanchen/VisRL.
@@ 1 Introduction
With the advent of large multimodal models, the intension in perception tasks can now be expressed in a highly flexible way, which is nature language. Common LMMs, such as LLaVA [42] and Qwen-VL [74], first encode visual signals into tokens, and then both visual and text tokens are jointly processed by the large language models to produce final outputs. Despite showing powerful ability in many tasks [89, 38, 14, 32, 75, 91], this kind of one-pass end-to-end LMMs still suffer from hallucinations [22] and do not explicitly address the intention-driven visual perception problem. Further extending this paradigm, instead of treating the entire process as a single black-box inference, recent works have proposed frameworks like Visual Chain-of-Thought (Visual CoT [61]). These methods introduce an explicit reasoning step where the model first predicts a bounding box representing the critical region to focus on, crops the image to extract this region, and then feeds the cropped visual input back to the multimodal model. By conditioning the final answer on both the query and the selected focus area, this approach not only improves interpretability but also helps leverage the multi-turn in-context learning capabilities of the underlying LMM.
@@ 2 Related Work
Recently, several methods have attempted to enhance models’ intention-driven visual perception capabilities. VisCoT [61] employs a multi-turn interpretable processing mechanism with bounding boxes that dynamically focus on visual inputs. Similarly, SpatialCoT [46] achieves spatial grounding through spatial coordinates, while SegLLM [76] leverages mask-labeled data to enable reasoning about complex user segmentation intentions. V* (SEAL) [79] provides an LLM-guided search mechanism for efficient visual querying. Besides, both MLLM-TPO [84] and VisionLLM v2 [78] achieve intention-driven perception by training decoders for specific downstream tasks. Additionally, MVoT [34]introduces a text-image-text reasoning paradigm by training on interleaved data. These methods generally follow the supervised learning (SL) paradigm, therefore heavily relying on dense-labeled data (e.g., bounding boxes, spatial coordinates, masks, multi-round reasoning conversations), which constrains their ability to scale further. Instead, VisRL uses rewards as learning supervisions and gets rid of the requirement on dense annotations. On the other hand, several methods employ tool-usage for enhancement. Specifically, AURORA [5] leverages specialized detection models, while Plug-and-Play [9] adopts a multi-agent framework. Besides, ViRReq [66] leverages the knowledge base to decompose visual recognition. However, these approaches rely on external models or knowledge, but not focusing on enhancing the intrinsic capabilities of the models themselves, while VisRL tries to completing the task via learning to reasoning by the model itself.
@@
Chain-of-Thought (CoT) reasoning plays an important role for LMMs. The methods can be broadly categorized into two types. (1) Text-thought methods [12, 68, 27, 17, 81] elicit textual CoT reasoning of LLMs in visual reasoning tasks by introducing text thinking tokens inspired by [23], guiding towards the final response. Differently, VisRL emphasizes that visual information should be involved in the reasoning process to fully leverage the strengths of multimodal models, rather than relying solely on language tokens for reasoning. (2) Multi-modal-thought methods involves multimodal information in the reasoning process. The Mind’s Eye [80] elicit spatial reasoning of LLMs by visualizing their reasoning traces. Additionally,  [46, 61, 76] first generate visual marks (e.g. bounding boxes, spatial coordinates, masks), and subsequently perform CoT reasoning based on these fine-grained visual marks. These works still uses SL for optimization while VisRL first explores RL in this direction. Besides, recent repositories [62, 47] also tried to enhance the bounding boxes (bboxes) generation through RL, but they uses ground truth bboxes to give rewards, which is still limited by dense annotation data.
@@ 3.1 Data Generation
Before RL training, we use SFT as a warm-up stage. This stage changes the original model $\mathcal{M}_{org}$ into $\mathcal{M}_{SFT}$, so $\mathcal{M}_{SFT}$ is $\mathcal{M}_{RL}^{0}$. This stage requires bounding box annotations, so we directly uses samples from VisCoT dataset [61]. Our intent here is just making the model have the basic ability to generate bounding boxes in specific format. Therefore, the used data amount is relatively small here (30k in ours and 438k in VisCoT). The second part of our source datasets consists of question-answer pairs from nine source datasets that span five distinct domains, with the majority of them being Visual Question Answering (VQA) and Image Captioning datasets. These are specifically used for constructing preference data for RL training, therefore requiring no bounding box annotations. We use additional 180k data here. Then, we explain how we construct preference data based on our source datasets in details.
@@
For the current question and image $(Q,I)$, if the win set and the loss set is not empty, this question and image and its corresponding generated data are preserved. The intuition here is that we apply a filter to the questions in the dataset, selecting those with a difficulty level suitable for the current model. Questions that are too difficult will prevent the model from generating meaningful answers, while questions that are too easy will result in the model producing correct answers across the board. Both of these types of questions will be filtered out during the current training round. However, as the model’s capabilities are updated and strengthened, less data are filtered in the next iteration, as indicated by Data Num. of VisRL-Full vs. VisRL-Full-Iter1 in Tab. 6.
@@ 4.5 Ablation Study
Data Generation Pipeline. Our data generation process follows a self-evolution paradigm and adopts an actor-critics framework, where the actor is the SFTed model and the critics refer to the pre-SFT model. In this study, we conduct ablation experiments on several modules within the pipeline. As shown in Tab. 6, our objective is to maximize the proportion of positive instances in the constructed preference dataset, specifically WP-LN (win-positive, lose-negative), while minimizing the proportion of negative instances, WN-LP (win-negative, lose-positive). Experimental results indicate that under our pipeline, the proportion of positive instances is 64.64%. After the first and second iterations – where the model undergoes full VisRL training and then participates in data construction again – this proportion increases by 3.18% and 2.30%, respectively. Beside, the number of valid data increases from 30k to 33k to 35k. This is because, through iterative training, the model’s capability improves to overcome bottlenecks where obtaining the correct answer was previously outside of its ability. Specifically, some questions are inherently too difficult for model previously, meaning that no matter how many times the model responds, the answer remains incorrect, making it impossible to construct win/lose preference.
@@ 5 Conclusion
In this paper, we propose VisRL, a framework for learning intention-driven visual perception abilities from task feedback. This approach enables the model to undergo RL through self-evolution when trained on simple data with only final responses. Specifically, we introduce a novel pipeline for generating CoT preference data based on the model’s own actor-critics process, eliminating the need for external models or human annotations. Using these self-constructed data, we further optimize visual perception in two RL stages: (1) independently optimizing generated bounding box; (2) jointly optimizing both generated bounding box and response. Extensive experiments on various benchmarks and LMMs demonstrate the effectiveness of the proposed framework, establishing a solid foundation for future exploration.
@@ Appendix B Visual Grounding
As shown in Tab. 8, VisRL surpasses all previous generalist models, even outperforming models with significantly larger parameter. Moreover, in most of the cases, our method exceeds the performance of previous state-of-the-art specialist models, e.g. G-DINO-L [44] and UNINEXT [83]. This demonstrates the exceptional capability of our approach in accurately predicting bounding boxes. Notably, our model achieves improvements of 1% to 5% to VisCoT. ”Top-1 Accuracy@0.5,” refers to the accuracy of a model in correctly predicting the bounding box as the top-ranked output when the IoU between the predicted and GT bounding boxes is at least 50%.

### v2

@@ Abstract
Visual understanding is inherently intention-driven—humans selectively focus on different regions of a scene based on their goals. Recent advances in large multimodal models (LMMs) enable flexible expression of such intentions through natural language, allowing queries to guide visual reasoning processes. Frameworks like Visual Chain-of-Thought have demonstrated the benefit of incorporating explicit reasoning steps, where the model predicts a focus region before answering a query. However, existing approaches rely heavily on supervised training with annotated intermediate bounding boxes, which severely limits scalability due to the combinatorial explosion of intention-region pairs. To overcome this limitation, we propose VisRL, the first framework that applies reinforcement learning (RL) to the problem of intention-driven visual perception. VisRL optimizes the entire visual reasoning process using only reward signals. By treating intermediate focus selection as an internal decision optimized through trial-and-error, our method eliminates the need for costly region annotations while aligning more closely with how humans learn to perceive the world. Extensive experiments across multiple benchmarks show that VisRL consistently outperforms strong baselines, demonstrating both its effectiveness and its strong generalization across different LMMs. Our code is available at https://github.com/zhangquanchen/VisRL.
@@ 1 Introduction
With the advent of large multimodal models, the intention in perception tasks can now be expressed in a highly flexible way, which is natural language. Common LMMs, such as LLaVA [43] and Qwen-VL [75], first encode visual signals into tokens, and then both visual and text tokens are jointly processed by the large language models to produce final outputs. Despite showing powerful ability in many tasks [90, 39, 14, 33, 76, 92, 15], this kind of one-pass end-to-end LMMs still suffer from hallucinations [23] and do not explicitly address the intention-driven visual perception problem. Further extending this paradigm, instead of treating the entire process as a single black-box inference, recent works have proposed frameworks like Visual Chain-of-Thought (Visual CoT [62]). These methods introduce an explicit reasoning step where the model first predicts a bounding box representing the critical region to focus on, crops the image to extract this region, and then feeds the cropped visual input back to the multimodal model. By conditioning the final answer on both the query and the selected focus area, this approach not only improves interpretability but also helps leverage the multi-turn in-context learning capabilities of the underlying LMM.
@@ 2 Related Work
Recently, several methods have attempted to enhance models’ intention-driven visual perception capabilities. VisCoT [62] employs a multi-turn interpretable processing mechanism with bounding boxes that dynamically focus on visual inputs. Similarly, SpatialCoT [47] achieves spatial grounding through spatial coordinates, while SegLLM [77] leverages mask-labeled data to enable reasoning about complex user segmentation intentions. V* (SEAL) [80] provides an LLM-guided search mechanism for efficient visual querying. Besides, both MLLM-TPO [85] and VisionLLM v2 [79] achieve intention-driven perception by training decoders for specific downstream tasks. Additionally, MVoT [35]introduces a text-image-text reasoning paradigm by training on interleaved data. These methods generally follow the supervised learning (SL) paradigm, therefore heavily relying on dense-labeled data (e.g., bounding boxes, spatial coordinates, masks, multi-round reasoning conversations), which constrains their ability to scale further. Instead, VisRL uses rewards as learning supervision and gets rid of the requirement on dense annotations. On the other hand, several methods employ tool-usage for enhancement. Specifically, AURORA [5] leverages specialized detection models, while Plug-and-Play [9] adopts a multi-agent framework. Besides, ViRReq [67] leverages the knowledge base to decompose visual recognition. However, these approaches rely on external models or knowledge, but not focusing on enhancing the intrinsic capabilities of the models themselves, while VisRL tries to completing the task via learning to reasoning by the model itself.
@@
Chain-of-Thought (CoT) reasoning plays an important role for LMMs. The methods can be broadly categorized into two types. (1) Text-thought methods [12, 69, 28, 18, 82] elicit textual CoT reasoning of LLMs in visual reasoning tasks by introducing text thinking tokens inspired by [24], guiding towards the final response. Differently, VisRL emphasizes that visual information should be involved in the reasoning process to fully leverage the strengths of multimodal models, rather than relying solely on language tokens for reasoning. (2) Multi-modal-thought methods involves multimodal information in the reasoning process. The Mind’s Eye [81] elicits spatial reasoning of LLMs by visualizing their reasoning traces. Additionally,  [47, 62, 77] first generate visual marks (e.g. bounding boxes, spatial coordinates, masks), and subsequently perform CoT reasoning based on these fine-grained visual marks. These works still uses SL for optimization while VisRL first explores RL in this direction. Besides, recent repositories [63, 48] also tried to enhance the bounding boxes (bboxes) generation through RL, but they uses ground truth bboxes to give rewards, which is still limited by dense annotation data.
@@ 3.1 Data Generation
Before RL training, we use SFT as a warm-up stage. This stage changes the original model $\mathcal{M}_{org}$ into $\mathcal{M}_{SFT}$, so $\mathcal{M}_{SFT}$ is $\mathcal{M}_{RL}^{0}$. This stage requires bounding box annotations, so we directly use samples from VisCoT dataset [62]. Our intent here is just to make the model have the basic ability to generate bounding boxes in specific format. Therefore, the used data amount is relatively small here (30k in ours and 438k in VisCoT). The second part of our source datasets consists of question-answer pairs from nine source datasets that span five distinct domains, with the majority of them being Visual Question Answering (VQA) and Image Captioning datasets. These are specifically used for constructing preference data for RL training, therefore requiring no bounding box annotations. We use an additional 180k data here. Then, we explain how we construct preference data based on our source datasets in details.
@@
For the current question and image $(Q,I)$, if the win set and the loss set are not empty, this question and image and its corresponding generated data are preserved. The intuition here is that we apply a filter to the questions in the dataset, selecting those with a difficulty level suitable for the current model. Questions that are too difficult will prevent the model from generating meaningful answers, while questions that are too easy will result in the model producing correct answers across the board. Both of these types of questions will be filtered out during the current training round. However, as the model’s capabilities are updated and strengthened, less data are filtered in the next iteration, as indicated by Data Num. of VisRL-Full vs. VisRL-Full-Iter1 in Tab. 6.
@@ 4.5 Ablation Study
Data Generation Pipeline. Our data generation process follows a self-evolution paradigm and adopts an actor-critics framework, where the actor is the SFTed model and the critics refer to the pre-SFT model. In this study, we conduct ablation experiments on several modules within the pipeline. As shown in Tab. 6, our objective is to maximize the proportion of positive instances in the constructed preference dataset, specifically WP-LN (win-positive, lose-negative), while minimizing the proportion of negative instances, WN-LP (win-negative, lose-positive). Experimental results indicate that under our pipeline, the proportion of positive instances is 64.64%. After the first and second iterations – where the model undergoes full VisRL training and then participates in data construction again – this proportion increases by 3.18% and 2.30%, respectively. Besides, the number of valid data increases from 30k to 33k to 35k. This is because, through iterative training, the model’s capability improves to overcome bottlenecks where obtaining the correct answer was previously outside of its ability. Specifically, some questions are inherently too difficult for model previously, meaning that no matter how many times the model responds, the answer remains incorrect, making it impossible to construct win/lose preference.
@@ 5 Conclusion
In this paper, we propose VisRL, a framework for learning intention-driven visual perception abilities from task feedback. This approach enables the model to undergo RL through self-evolution when trained on simple data with only final responses. Specifically, we introduce a novel pipeline for generating CoT preference data based on the model’s own actor-critic process, eliminating the need for external models or human annotations. Using these self-constructed data, we further optimize visual perception in two RL stages: (1) independently optimizing generated bounding box; (2) jointly optimizing both generated bounding box and response. Extensive experiments on various benchmarks and LMMs demonstrate the effectiveness of the proposed framework, establishing a solid foundation for future exploration.
@@ Appendix B Visual Grounding
As shown in Tab. 8, VisRL surpasses all previous generalist models, even outperforming models with significantly larger parameters. Moreover, in most of cases, our method exceeds the performance of previous state-of-the-art specialist models, e.g. G-DINO-L [45] and UNINEXT [84]. This demonstrates the exceptional capability of our approach in accurately predicting bounding boxes. Notably, our model achieves improvements of 1% to 5% to VisCoT. ”Top-1 Accuracy@0.5,” refers to the accuracy of a model in correctly predicting the bounding box as the top-ranked output when the IoU between the predicted and GT bounding boxes is at least 50%.